Operation Summary for August 26, 2011 - August 29, 2011 |
Wednesday August 31 The day shift began with Operations monitoring the Tevatron (TEV) store 9048 (139.53E30) and the Recycler stash (97.2E10), with the Antiproton Source (Pbar) stacking (20.6E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, at 10:55 AM, operators reported problems with a Pbar kicker (D:EKIK); kicker personnel investigated and reported that its high voltage cables were arcing again. They re-terminated the cables. Operators disabled the Booster beam for about 50 minutes to open up the West Anode Power Supply to look for LCW leaks; there were no leaks. Booster personnel tuned the Booster beam to run without two RF stations (BRF13 & BRF15). At 1:16 PM, the TEV suffered a beam abort due to a bad vacuum indication at TEV sector B4; operators investigated and reported that the vacuum was normal. Booster RF personnel brought BRF13 online at 2:17 PM. Operations put the TEV into shot setup at 2:34 PM, but at 3:26 PM, the TEV suffered a sector DA quench during the seventh antiproton injection due to a kicker failure (T:E48K). During the evening shift, Operations established beam to the Meson FTBFT-1015experiment at 4:10 PM. Operators put the TEV through a dry squeeze at 5:02 PM. A TEV RF station (TRF4) tripped off during a wet squeeze at 5:34 PM; a RF person reset the station. Operations put the TEV into shot setup at 7:27 PM. Operators reported at 7:34 PM that they received an alarm from a dehumidifier holding tank at the MI-12 service building; which will be investigated later. At 8 PM, operators took a look in the TEV tunnel via a TEV sector F0 camera. There was a lot of water on the floor. Operators contacted MiniBooNE, Safety personnel, and others. An RF person reported that there was a known small leak in a RF Cavity that must have gotten worse; he came in to investigate further. Operators put an energy conservation time line into the Time Line Generator (TLG) at 9:16 PM. Operators prepared F sector for an access. The RF person reported finding a leak on top of cavity TRF5; he had the leak fixed by 10:24 PM. Operations reestablished beam to MiniBooNE, NuMI, and Pbar by 11:29 PM. Thursday September 1 The midnight shift began with the TEV turn on in progress, with Operations monitoring the stash (208.4E10), with Pbar stacking (9.4E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, operators put the TEV through a wet squeeze at 1:32 AM. Operations put the TEV into shot setup at 1:55 AM and established store 9053 at 2:54 AM with an average initial luminosity (AIL) of 250.5E30. Neutron Therapy Facility (NTF) personnel began tuning, at 7:40 AM, in preparation for treating two patients later in the morning. During the day shift, at 9:28 AM, BRF4 tripped off for the first of three times in 16 minutes; RF personnel bypassed the station and made repairs. Operations halted beam, at 12:28 PM, in preparation for the required ComEd power curtailment test. Operators prepared the Booster, MI, and Pelletron for accesses. Operators prepared the I- Source and the 8 GeV beamline for accesses at 1:34 PM. TEV personnel began crystal collimator studies at 2:27 PM. Operations reported at 3:06 PM that the power curtailment test for ComEd had ended. Operators reported at 3:50 PM that all the machines that were in access had come out of access. Operations resumed sending beam to NuMI, MiniBooNE, and Pbar. During the evening shift, Operations established beam to T-1015 at 4:43 PM. Operations terminated store 9053 at 6:14 PM. At 6:37 PM, BRF8 tripped off for the first of three times in about two and a half hours; it reset. At 6:51 PM, BRF7 tripped off for the first of five times in about four hours; it reset. Operations established store 9054 at 7:33 PM with an AIL of 275.49E30. Operations disabled beam to T-1015 at 9 PM. Friday September 2 The midnight shift began with Operations monitoring store 9054 (131.2E30) and the stash (140.1E10), with Pbar stacking (11.7E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, at 4:41 AM, operators reported that six TEV sector F0 ion pumps tripped off; they reset, but operators waited for another ion pump to recover the vacuum before turning the others back on. Also, this held off stacking until 5:14 AM. NTF began tuning, at 7:25 AM, in preparation to treat a single patient later in the morning. The Plans for Friday and the Weekend The plans for today and the weekend are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 09/02/2011 email Fermilab |
| Security, Privacy, Legal | |